OpenStudy (anonymous):

at 1:00 pm a thermometer reading 10F is removed from a freezer and placed in a room whose temperature is 65 F. at 1:05 pm, the thermometer reads 25 F. Later, the thermometer is placed back in the freezer. at 1:30 pm the thermometer reads 32 F. when was the thermometer returned to the freezer and what was the thermometer reading at that time? Answer: 1:20 pm; 50.15 F

OpenStudy (anonymous):

So you answered your own question. makes the job much easier for us now doesn't it? :D

OpenStudy (anonymous):

That looks like a physics question. We might need some associated thermal formulas. I have not had a heat ransfer course in years so I am a little rusty in this field. What course is this question for?

OpenStudy (anonymous):

looks like some sort of newton's law of cooling puzzle

OpenStudy (anonymous):

There is an answer given to us but i don't know the solution. im stuck at the point where the temp is heated and cooled after.

OpenStudy (anonymous):

It is a newtons law of cooling problem

OpenStudy (anonymous):

@znerte Are you studying law of exponential change?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so far ive got the k for this question which is -.0637

OpenStudy (anonymous):

lets see the temperature difference starts at 55 degrees at 1:00 five minutes later it is 40 degrees, so you can model this with \[A(t)=55\left(\frac{8}{11}\right)^{\frac{t}{5}}\]

OpenStudy (anonymous):

or you can also do it by using the (rather slavish) formula \[55e^{whatever}\]

OpenStudy (anonymous):

now i am kind of stuck, but i guess we are to assume that the freezer is a constant temperature of ten degrees, right?

OpenStudy (anonymous):

Good assumption @satellite73

OpenStudy (anonymous):

i think so because it is at 10F when it is removed from the freezer

OpenStudy (anonymous):

lol i am a bit slow this morning so far all i have is the temperature of the thermometer outside of the room is \[A(t)=65-55e^{-.067t}\]

OpenStudy (anonymous):

i meant outside of the freezer

OpenStudy (anonymous):

Basically we apply Newton's Law of Cooling as mentioned earlier. We can derive the formula using separable DEs/law of exponential change but I will choose not to do that. Going directly to the problem, we know that the initial temperature inside the freezer at 1 pm was 10F. It was then placed inside the room which was at temperature 65F and at 1:05 pm, 5min later, the thermometer reads 25F. From this data here and using Newton's Law of Cooling, we get:\[\bf T-T_s=(T_0-T_s)e^{-kt} \rightarrow 25-65=(10-65)e^{-k(5)}\]\[\bf \implies -40=(-55)e^{-5k} \implies e^{-5k}=\frac{ 8 }{ 11 }\]\[\bf \implies -5k = \ln \left( \frac{ 8 }{ 11 } \right) \implies k=\ln \left( \frac{ 11 }{ 8 } \right)^{\frac{1}{5}} \approx -0.0637\] Good so far? @znerte

OpenStudy (anonymous):

he/she wrote that above

OpenStudy (anonymous):

I shall move on then.

OpenStudy (anonymous):

yeah please, because i can figure out the rate of warming, but not the rate of cooling want to see how you do it

OpenStudy (anonymous):

question is the k is the same at the rate of cooling

OpenStudy (anonymous):

why would it be?

OpenStudy (anonymous):

ok lets assume that it is the same what equation can we solve?

OpenStudy (anonymous):

it is out for \(t\) minutes and back in for \(30-t\) minutes

OpenStudy (anonymous):

holy crap i think i got it!

OpenStudy (anonymous):

lets see if this makes sense we assume the rate of cooling is the same as the rate of heating since i know literally nothing about physics lets just assume it for a second okay?

OpenStudy (anonymous):

you have found the rate at \(-.0637\) so the equation for the temperature of the thermometer outside the freezer is \[65-55e^{-.0637t}\]

OpenStudy (anonymous):

at some time \(t_0\) it is put in the freezer and the temperature is \[65-55e^{-.0637t_0}\] i just used \(t_0\) to indicate that the above is some constant that we do not know now assuming the rate hasn't changed, the temperature once it is back in the freezer is \[\left(65-55e^{-.0637t_0}\right)e^{-.0637t}+10\]

OpenStudy (anonymous):

out for \(t_0\) minutes, then back in for \(30-t_0\) minutes

OpenStudy (anonymous):

\[\ln \frac{ 32-10 }{ 65-55^{-.0637t}-10 }=-0.0637(30)+0.0637t\]

OpenStudy (anonymous):

lets try solving \[\left(65-55e^{-.0637t_0}\right)e^{-.0637(30-t_0)}+10=32\]

OpenStudy (anonymous):

solving for t i get 20.555

OpenStudy (anonymous):

then go with that, because my answer is not right !